OpenStudy (anonymous):

Find a number Z so that P{ N(4, 4) < Z } = .75 ( the probability that a normal 4, 4 is less than Z equals .75) can you show steps to see how you did it thanks

OpenStudy (zarkon):

is your notation N(\(\mu,\sigma^2\)) or N(\(\mu,\sigma\))

OpenStudy (anonymous):

2nd one

OpenStudy (zarkon):

so the mean is 4 and the standard deviation is 4

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

your notation for the problem is not very nice...I'm going to modify it a little and you tell me if it is kn to you Let \(X\) have a normal distribution with mean 4 and standard deviation 4 you wint to find the value of \(x\) such that \[P(X\le x)=.75\]

OpenStudy (zarkon):

*want to find...

OpenStudy (anonymous):

um i dont even know how to start it

OpenStudy (zarkon):

are you ok with how I worded the problem?

OpenStudy (anonymous):

so the z score is .7734

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

where do you get that?

OpenStudy (zarkon):

I see what you did...no that is not correct

OpenStudy (zarkon):

you need to go backwards from what you just did

OpenStudy (anonymous):

what do you mean backwards

OpenStudy (zarkon):

\[P(X\le x)=.75\] \[\Rightarrow P\left(\frac{X-4}{4}\le \frac{x-4}{4}\right)=.75\] \[\Rightarrow P\left(Z\le \frac{x-4}{4}\right)=.75\] \[\Rightarrow P\left(Z\le z\right)=.75\] looking up on a table (or a good calculator) one should get z=.6745

OpenStudy (zarkon):

my calculator says it is z=.67448974945

OpenStudy (zarkon):

but \[z=\frac{x-4}{4}\] solve for x gives \[x=4z+4\] plugging in the value of z from above and we get \[x=6.6979589978\]

OpenStudy (anonymous):

how do you figure the z cause when i look for 0.75 i get o.7734

OpenStudy (zarkon):

you are looking it up wrong...the .75 is a probability ...look in the middle of the table for a number that is closest to .75

OpenStudy (anonymous):

oh that makes sense now k thanks

OpenStudy (zarkon):

good